OpenStudy (anonymous):

Solve each equation 16. 3x-12=2x 17. -7x+7=3x-23 18. 0.6x-1.8=1.2x

OpenStudy (phi):

for 16. write -2x on each side (add -2x to both sides) then simplify

OpenStudy (anonymous):

okay then divide

OpenStudy (anonymous):

-12?

OpenStudy (phi):

3x-2x-12=2x-2x

OpenStudy (phi):

you should get in the habit of writing the entire equation. Less chance of making a mistake

OpenStudy (anonymous):

i did on paper just not on here

OpenStudy (phi):

3x-2x-12=2x-2x 3x-2x is 1x (3 x's take away 2 x's leaves 1 x) and of course 2x-2x is 0 so x -12 =0 now what do you do?

OpenStudy (anonymous):

can it be 1x-12=0? expect for x-12=0?

OpenStudy (phi):

1x is the same as x. (math people don't bother writing 1 ) anyway x-12= 0 what do you do next to find x?

OpenStudy (anonymous):

+12 on both sides

OpenStudy (phi):

yes

OpenStudy (anonymous):

x=12

OpenStudy (phi):

17. -7x+7=3x-23 you could add +7x to both sides and add +23 to both sides or, add -3x to both sides and -7 to both sides.

OpenStudy (anonymous):

to both sides of what?

OpenStudy (phi):

when solving an equation you do the same thing to both sides (to be fair) both sides means both sides of the = sign

OpenStudy (anonymous):

im doing the problem right now

OpenStudy (anonymous):

x=3

OpenStudy (phi):

17. -7x+7=3x-23 add +7x to both sides. write +7x on each side 7x-7x+7= 3x+7x-23 7x-7x on the left is 0 3x+7x on the right is 10x so 7= 10x-23 add 23 to both sides 23+7= 10x-23+23 30= 10x now what do you do?

OpenStudy (anonymous):

got the answer already 3.

OpenStudy (phi):

yes. 10x= 30 (changed from 30=10x from habit) divide by 10 on both sides x= 30/10 x=3

OpenStudy (anonymous):

okay #18

OpenStudy (phi):

18. 0.6x-1.8=1.2x what would you do first?

OpenStudy (anonymous):

+0.6 on both sides

OpenStudy (phi):

try -0.6x

OpenStudy (anonymous):

or subtract

OpenStudy (anonymous):

okay doing the problem noww

OpenStudy (anonymous):

is it -1.8/0.6

OpenStudy (anonymous):

x=-3

OpenStudy (phi):

x= -1.8/0.6 yes to make it (look) easy, multiply top and bottom by 10 \[ x= -\frac{1.8}{0.6}\cdot \frac{10}{10} \] (10 over 10 is one, so this does not change the value)

OpenStudy (anonymous):

how'd you'd get 10?

OpenStudy (phi):

10 times 0.6 = 6 (no decimal) 10 * 1.8= 18 so the problem becomes - 18/6 which is easy you get -3

OpenStudy (anonymous):

OH WELL i did it a diffrent way and still got -3

OpenStudy (phi):

how did you do it ?

OpenStudy (anonymous):

after i subtracted 0.6 on both sides. i got 0.6... i brought down the -1.8 and divide it by 0.6 i got -3 for x

OpenStudy (phi):

how did you divide it? with a calculator ?

OpenStudy (anonymous):

yeah.

OpenStudy (phi):

if you want to do it without a calculator, you use the trick I showed you. (assuming you know that 3*6=18!)

OpenStudy (anonymous):

ohhh thats pretty easy

OpenStudy (phi):

You are smarter than the average bear.

OpenStudy (anonymous):

lol thanks for the help